### v1

# sample.bib

Abstract
Machine unlearning algorithms aim to efficiently remove data from a model without retraining it from scratch, in order to enforce data privacy, remove corrupted or outdated data, or respect a user’s “right to be forgotten.” Certified machine unlearning is a strong theoretical guarantee that quantifies the extent to which data is erased from the model weights. Most prior works in certified unlearning focus on models trained on convex or strongly convex loss functions, which benefit from convenient convergence guarantees and the existence of global minima. For nonconvex objectives, existing algorithms rely on limiting assumptions and expensive computations that hinder practical implementations. In this work, we propose a simple first-order algorithm for unlearning on general nonconvex loss functions which unlearns by “rewinding” to an earlier step during the learning process and then performs gradient descent on the loss function of the retained data points. Our algorithm is black-box, in that it can be directly applied to models pretrained with vanilla gradient descent with no prior consideration of unlearning. We prove $(\epsilon,\delta)$ certified unlearning and performance guarantees that establish the privacy-utility-complexity tradeoff of our algorithm, with special consideration for nonconvex functions that satisfy the Polyak-Lojasiewicz inequality.
1 Introduction
Machine unlearning, or data deletion from models, refers to the problem of removing the influence of some data from a trained model without the computational expenses of completely retraining it from scratch [caoOGMUL]. This research direction has become highly relevant in the last few years, due to increasing concern about user privacy and data security as well as the growing cost of retraining massive deep learning models on constantly updated datasets. For example, recent legislation that protects a user’s “Right to be Forgotten,” including the European Union’s General Data Protection Regulation (GDPR), the California Consumer Privacy Act (CCPA), and the Canadian Consumer Privacy Protection Act (CPPA), mandate that users be allowed to request removal of their personal data, which may be stored in databases or memorized by models [sekharirememberwhatyouwant]. In addition, machine unlearning has practical implications for removing the influence of corrupted, outdated, or mislabeled data [nguyen2022surveymachineunlearning] [kurmanji2023towards].
The general goal of machine unlearning algorithms is to yield a model that resembles the model obtained from a full retrain on the updated dataset after data is removed. This requirement is formalized in the concept of “certified unlearning,” a strong theoretical guarantee inspired by differential privacy that probabilistically bounds the difference between the model weights returned by the unlearning and retraining algorithms [Guo]. However, an algorithm only satisfying certified unlearning is not sufficient, as many such algorithms may still be impractical to use. For example, a trivial unlearning algorithm retrains the model from scratch on the retained dataset, but this would provide no efficiency gain. On the other hand, unlearning is also provably satisfied if the learning and unlearning algorithm both output weights randomly sampled from the same Gaussian distribution, but this would yield a poorly performing model. Therefore, an ideal algorithm optimally balances data deletion, model accuracy, and computation, also known as the “privacy-utility-efficiency” tradeoff. Moreover, because training from scratch is computationally expensive in the machine unlearning setting, we desire black-box unlearning algorithms, which can be applied to pretrained models and do not require training to complete with the intention of unlearning later.
Most certified unlearning algorithms are designed for convex or strongly convex functions [Guo] [neel21a] [sekharirememberwhatyouwant]. Relaxing the convexity requirement is challenging since nonconvex functions do not have unique global minima. Recently, there have been several works that provide certified unlearning guarantees for nonconvex functions. For example, [zhang2024towards] proposes a single-step Newton update algorithm with convex regularization followed by Gaussian perturbation, inspired by existing second-order unlearning methods such as [Guo] [sekharirememberwhatyouwant]. Similarly, [qiao2024efficientgeneralizablecertifiedunlearning] proposes a quasi-second-order method that exploits Hessian vector products to avoid directly computing the Hessian. Finally, [chien2024langevin] proposes a first-order method in the form of projected noisy gradient descent with Gaussian noise added at every time step. These algorithms are listed in Table 1.
Prior work has focused on achieving theoretical unlearning guarantees in the nonconvex setting; however, practical unlearning algorithms should also be computationally efficient, convenient to use, unbounded by restrictive assumptions, and yield models that perform well. First, [zhang2024towards] and [qiao2024efficientgeneralizablecertifiedunlearning] are both (quasi) second-order methods, but first-order methods are more computationally efficient. Second, [chien2024langevin] [qiao2024efficientgeneralizablecertifiedunlearning] are not black-box, since during the training process, [chien2024langevin] requires injecting Gaussian noise at every step and [qiao2024efficientgeneralizablecertifiedunlearning] requires recording a statistical vector for each data sample at each time step. These “white-box” algorithms require significant changes to standard learning algorithms, which hinders practical implementation. Finally, both [zhang2024towards] and [chien2024langevin] require a bounded feasible parameter set, which they leverage to provide a loose bound on the distance between the retraining and unlearning outputs dependent on the parameter set radius $R$. However, this yields excessive noise requirements, with the standard deviation $\sigma$ scaling linearly or polynomially with $R$, translating to degraded model performance. See Table 1 for a summary of these results and Section 3 for a more in-depth discussion of $R$ and comparison with prior literature.
Algorithm Loss Function Method Black-box? Bounded $||\theta||\leq R?$
Newton Step [Guo] Strongly convex Second-order No No
Descent-to-delete (D2D) [neel21a] Strongly convex First-order Yes No
Langevin unlearning [chien2024langevin] Nonconvex First-order No Yes
Constrained Newton Step [zhang2024towards] Nonconvex Second-order Yes Yes
Hessian-free Unlearning [qiao2024efficientgeneralizablecertifiedunlearning] Nonconvex Quasi-second-order No No
Our Work (R2D) Nonconvex First-order Yes No
Table 1: Comparison of certified unlearning algorithms for convex and nonconvex functions. The model parameters are denoted by $\theta$.
In this work, we introduce “rewind-to-delete” (R2D), a first-order, black-box, certified unlearning algorithm for general nonconvex functions that does not depend on the parameter set radius $R$. Our learning algorithm consists of vanilla gradient descent steps on the loss function of the dataset followed by Gaussian perturbation. To remove some data from the model, our unlearning algorithm “rewinds” to a model checkpoint from earlier in the original training trajectory, performs additional gradient descent steps on the new loss function, and again adds Gaussian perturbation to the weights. The checkpoint can be saved during the training period, or it can be computed post hoc from a pretrained model via the proximal point method. We show that this algorithm yields $(\epsilon,\delta)$ certified unlearning and provide theoretical guarantees that explicitly address the “privacy-utility-efficiency” tradeoff [liutrilemma] of our unlearning algorithm. Our analysis relies on simple assumptions and does not require a bounded parameter set or projection. In addition, the algorithm is black-box, as it can be applied to a pretrained model as long as the model was trained with vanilla gradient descent.
We also analyze the special case of nonconvex loss functions that satisfy the Polyak-Łojasiewicz (PL) inequality. The PL inequality is a weak condition that guarantees the existence of a connected basin of global minima, to which gradient descent converges at a linear rate [karimiPL]. These functions are especially relevant to deep learning due to the recent discovery that overparameterized neural networks locally satisfy the PL condition in neighborhoods around initialization [LIUPLoverparametrized].
Our contributions are as follows.
  • We develop a novel ($\epsilon$, $\delta$) certified unlearning algorithm for general nonconvex loss functions. To the best of our knowledge, this is the first nonconvex certified unlearning algorithm that is first-order and black-box.
  • We provide theoretical unlearning guarantees that demonstrate the privacy-utility-efficiency tradeoff of our algorithm, allowing for controllable noise at the expense of privacy or computation. Our analysis requires significantly fewer assumptions than prior work; notably, we do not require that the parameter set is uniformly bounded.
  • We provide utility guarantees of our algorithm, and we show that for the special case of PL loss functions, we obtain linear convergence guarantees and generalization guarantees.
1.1 Related Work
Differential privacy.
Differential privacy [dwork2006] is a well-established framework designed to protect individual privacy by ensuring that the inclusion or exclusion of any single data point in a dataset does not significantly affect the output of data analysis or modeling, limiting information leakage about any individual within the dataset. Specifically, a differentially private learning algorithm yields a model trained on some dataset that is probabilistically indistinguishable from a model trained on the same dataset with a data sample removed or replaced. The concept of $(\epsilon,\delta)$-privacy quantifies the strength of this privacy guarantee in terms of the privacy loss, $\epsilon$, and the probability of a privacy breach, $\delta$ [dwork2014algorithmic]. This is typically achieved during or after training by adding controlled noise to the data, model weights [wublackboxdp] [zhangDPijcai2017p548], gradients [abadiDP] [zhangDPijcai2017p548], or objective function [chaudhuri11aDP], in order to mask information about any one sample in the dataset. However, greater noise typically corresponds to worse model performance, leading to a trade-off between utility and privacy. The theory and techniques of differential privacy provide a natural starting point for the rigorous analysis of unlearning algorithms.
As observed in [wublackboxdp], “white-box” differentially private algorithms, which require code changes at every training step to inject noise, are challenging to deploy in the real world due to additional development and runtime overhead incurred from implementing a nonstandard learning algorithm. Rather than adding noise at each iteration, [wublackboxdp] and [zhangDPijcai2017p548] propose only perturbing the output of vanilla stochastic gradient descent, an approach which is easier to integrate into realistic development pipelines. In similar fashion, our proposed black-box unlearning algorithm can be implemented without any special steps during learning with gradient descent, and we also mask the difference in algorithm outputs with noise at the end of training. The difference between our approach and [wublackboxdp] [zhangDPijcai2017p548], however, is that our approach can accommodate the nonconvex case, leveraging model checkpointing to control the distance from the retraining trajectory.
Certified unlearning.
The term “machine unlearning” was first coined by [caoOGMUL] to describe a deterministic data deletion algorithm, which has limited application to general optimization problems. In the following years, techniques have been developed for “exact unlearning,” which exactly removes the influence of data, and “approximate unlearning,” which yields a model that is approximately close to the retrained model with some determined precision [xusurvey]. Our work focuses on the latter. Both [ginartmakingAI] and [Guo] propose a probabilistic notion of approximate unlearning, where the unlearning and retraining outputs must be close in distribution. Inspired by differential privacy, [Guo] introduces the definition of certified $(\epsilon,\delta)$ unlearning that we use in this work. Like differentially private algorithms, certified unlearning algorithms typically require injected noise to the weights or objective function, which can degrade model performance. However, unlearning algorithms are also designed to reduce computation, leading to a three-way trade-off between privacy, utility, and complexity.
Certified unlearning has been studied for a variety of specific settings, including linear and logistic models [Guo] [izzo21a], graph neural networks [chien2022certifiedgraphunlearning], and minimax models [liu2023certified], as well as convex models [sekharirememberwhatyouwant] [neel21a] [chien2024stochastic] [suriyakumar2022algorithms] and nonconvex models [chien2024langevin] [qiao2024efficientgeneralizablecertifiedunlearning] [zhang2024towards]. Specifically, our work is closely inspired by [neel21a], which proposes “descent-to-delete,” a first-order unlearning algorithm for strongly convex functions that unlearns by fine-tuning with gradient descent iterates on the loss function of the retained samples. The algorithms for general models can be categorized as first-order methods that only require access to the function gradients [neel21a] [chien2024stochastic] or second-order methods that leverage information from the Hessian to approximate the model weights that would result from retraining [sekharirememberwhatyouwant] [suriyakumar2022algorithms] [zhang2024towards] [qiao2024efficientgeneralizablecertifiedunlearning].
Nonconvex unlearning.
In addition to the works mentioned above, there are many approximate unlearning algorithms designed for nonconvex functions such as neural networks that have weaker theoretical guarantees. For example, [bui2024newtonsmethodunlearnneural] proposes a “weak unlearning” algorithm, which considers indistinguishability with respect to model output space instead of model parameter space. This is a weaker guarantee because the model weights may still retain information about the unlearned data and be susceptible to membership inference attacks. Another popular algorithm for neural networks is SCRUB [kurmanji2023towards], a gradient-based algorithm that balances maximizing error on the forget set and maintaining performance on the retain set. An extension of SCRUB, SCRUB+Rewind, “rewinds” the algorithm to a point where the error on the forget set is “just high enough,” so as to impede membership inference attacks. Finally, although [golatkar1] [golatkar2] propose unlearning algorithms for deep neural networks, they only provide a general upper bound on the amount of information retained in the weights rather than a strict certified unlearning guarantee.
Additional approaches include subtracting out the impact of the forget datapoints in each batch of gradient descent [Graves_Nagisetty_Ganesh_2021], gradient ascent on the loss function of the forget set [jang-etal-2023-knowledge], and retraining the last layers of the neural network on the retain set [goel]. Ultimately, while the ideas of checkpointing, gradient ascent, and “rewinding” have been considered in other machine unlearning works, our algorithm combines these elements in a novel manner to obtain strong theoretical guarantees that prior algorithms lack.
2 Algorithm
Let $\mathcal{D}=\{z_{1},...,z_{n}\}$ be a training dataset of $n$ data points drawn independently and identically distributed from the sample space $\mathcal{Z}$, and let $\Theta$ be the model parameter space. Let $A:\mathcal{Z}^{n}\to\Theta$ be a (randomized) learning algorithm that trains on $\mathcal{D}$ and outputs a model with weight parameters $\theta\in\Theta$, where $\Theta$ is the (potentially infinite) space of possible model weights. Typically, the goal of a learning algorithm is to minimize $f_{\mathcal{D}}(\theta)$, defined as the empirical loss on $\mathcal{D}$ as follows
$f_{\mathcal{D}}(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{z_{i}}(\theta)$
where $f_{z_{i}}(\theta)$ represents the loss on the sample $z_{i}$.
Let us “unlearn” or remove the influence of a subset of data $Z\subset\mathcal{D}$ from the output of the learning algorithm $A(\mathcal{D})$. Let $\mathcal{D^{\prime}}=\mathcal{D}\backslash Z$, and we denote by $U(A(\mathcal{D}),\mathcal{D},Z)$ the output of an unlearning algorithm $U$. The goal of the unlearning algorithm is to output a model parameter that is probabilistically indistinguishable from the output of $A(\mathcal{D^{\prime}})$. This is formalized in the concept of $(\epsilon,\delta)$-indistinguishability, which is used in the differential privacy literature to characterize the influence of a data point on the model output [dwork2014algorithmic].
Definition 2.1.
[dwork2014algorithmic] [neel21a] Let $X$ and $Y$ be random variables over some domain $\Omega$. We say $X$ and $Y$ are $(\epsilon,\delta)$-indistinguishable if for all $S\subseteq\Omega$ we have
$\mathbb{P}[X\in S]\leq e^{\epsilon}\mathbb{P}[Y\in S]+\delta,$
$\mathbb{P}[Y\in S]\leq e^{\epsilon}\mathbb{P}[X\in S]+\delta.$
In the context of differential privacy, $X$ and $Y$ are the learning algorithm outputs on neighboring datasets that differ in a single sample. $\epsilon$ is the privacy loss or budget, which can be interpreted as a limit on the amount of information about an individual that can be extracted from the model, whereas $\delta$ accounts for the probability that these privacy guarantees might be violated. Definition 2.1 extends naturally to the following definition of $(\epsilon,\delta)$ certified unlearning.
Definition 2.2.
Let $Z$ denote a subset of the training dataset which we would like to unlearn. Then $U$ is an $(\epsilon,\delta)$ certified unlearning algorithm for $A$ if for all such $Z$, $U(A(\mathcal{D}),\mathcal{D},Z)$ and $A(\mathcal{D}\backslash Z)$ are $(\epsilon,\delta)$-indistinguishable.
Next, we describe our algorithms for machine unlearning. The learning algorithm $A$ (Algorithm 1) performs gradient descent updates on $f_{\mathcal{D}}$ for $T$ iterations, and the iterate at the $T-K$ time step is saved as a checkpoint or computed post hoc via the proximal point algorithm (Algorithm 3). Then Gaussian noise is added to the final parameter $\theta_{T}$, and the perturbed parameter is used for model inference. When a request is received to remove the data subset $Z$, the checkpointed model parameter $\theta_{T-K}$ is loaded as the initial point of the unlearning algorithm $U$ (Algorithm 2). Then we perform $K$ gradient descent steps on the new loss function $f_{\mathcal{D^{\prime}}}$ and add Gaussian noise again to the final parameter, using the perturbed weights for future model inference.
Algorithm 1 $A$: Rewind-To-Descent (R2D) Learning Algorithm
dataset $\mathcal{D}$, initial point $\theta_{0}\in\Theta$
for t = 1, 2, …, T do
     $\theta_{t}=\theta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})$triangleright Gradient descent steps
end for
Save checkpoint $\theta_{T-K}$ or compute $\theta_{T-K}$ via Algorithm 3
Sample $\xi\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{d})$
$\tilde{\theta}=\theta_{T}+\xi$triangleright Gaussian perturbation
Use $\tilde{\theta}$ for model inference
Upon receiving an unlearning request, execute Algorithm 2
Algorithm 2 $U$: R2D Unlearning Algorithm
dataset $\mathcal{D^{\prime}}$, model checkpoint $\theta_{T-K}$
$\theta^{\prime\prime}_{0}=\theta_{T-K}$
for t = 1, …, K do
     $\theta^{\prime\prime}_{t}=\theta^{\prime\prime}_{t-1}-\eta\nabla f_{\mathcal{D}^{\prime}}(\theta^{\prime\prime}_{t-1})$triangleright Gradient descent steps
end for
Sample $\xi\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{d})$
$\tilde{\theta}^{\prime\prime}=\theta^{\prime\prime}_{K}+\xi$triangleright Gaussian perturbation
Use $\tilde{\theta}^{\prime\prime}$ for model inference
When a model is trained without a checkpoint saved, we can still obtain a black-box unlearning algorithm by carefully “rewinding” the gradient descent training steps via the proximal point method, outlined in Algorithm 3. In standard gradient descent, we obtain $\theta_{t+1}$ from $\theta_{t}$ as follows,
$\theta_{t+1}=\theta_{t}-\eta\nabla f_{\mathcal{D}}(\theta_{t}).$
When we solve for previous gradient descent iterates, we obtain the following implicit equation:
$\theta_{t}=\theta_{t+1}+\eta\nabla f_{\mathcal{D}}(\theta_{t}).$ (1)
We note that the backward Euler update in (1) is distinct from a standard gradient ascent step, which is a forward Euler method. Instead, we can compute $\theta_{t}$ from $\theta_{t+1}$ by taking advantage of the connection between backward Euler for gradient flow and the proximal point method, an iterative algorithm for minimizing a convex function [Martinet_1970prox]. Let $g(\theta)$ denote a convex function. The proximal point method minimizes $g(\theta)$ by taking the proximal operator with parameter $\gamma$ of the previous iterate, defined as follows
$\theta_{k+1}=prox_{g,\gamma}(\theta_{k})=\arg\min_{x}\{g(x)+\frac{1}{2\gamma}||x-\theta_{k}||^{2}\}.$
Although for our problem, $f$ is nonconvex, sufficient regularization can produce a convex and globally tractable proximal point subproblem.
Lemma 2.3.
[drusvyatskiy2017proximalpointmethodrevisited] If $f(\theta)$ is continuously differentiable with $L$-Lipschitz gradient, $-f(\theta)+\frac{L}{2}||\theta||^{2}$ is convex.
Finally, by computing the proximal operator with respect to $-f(\theta)$, we can solve the implicit gradient ascent step.
Lemma 2.4.
Suppose $f(\theta)$ is continuously differentiable, $\theta_{t}$ is defined as in (1), and let $\eta<\frac{1}{L}$. Then $\theta_{t}=prox_{-f,\eta}(\theta_{t+1})$.
Proof.
Suppose $\theta^{*}=prox_{-f,\eta}(\theta_{t+1})=\arg\min_{x}\{-f(x)+\frac{1}{2\eta}||x-\theta_{t+1}||^{2}\}$, which is well-defined due to Lemma 2.3. Then the gradient of the objective function is zero at $\theta^{*}$ and thus
$-\nabla f(\theta^{*})+\frac{1}{\eta}(\theta^{*}-\theta_{t+1})=0,$
$\theta^{*}=\theta_{t+1}+\eta\nabla f(\theta^{*}).$
Algorithm 3 Compute Checkpoint via Proximal Point Method
datasets $\mathcal{D}$, model checkpoint $\theta_{T}$
$\theta^{\prime\prime}_{0}=\theta_{T}$
for t = 1, …, K do
     $\theta^{\prime\prime}_{t}=\arg\min_{x}\{-f_{\mathcal{D}}(x)+\frac{1}{2\eta}||x-\theta^{\prime\prime}_{t-1}||^{2}\}$triangleright Proximal point subproblem
end for
return $\theta^{\prime\prime}_{K}$
If $\eta<\frac{1}{L}$, then due to convexity, we can solve the proximal point subproblem easily via gradient descent or Newton’s method. Computationally, when $K\ll T$, the algorithm is comparable to other second-order unlearning methods that require a single Newton step. In addition, we only need to compute the model checkpoint once prior to unlearning requests, so this computation can be considered “offline” [izzo21a] [qiao2024efficientgeneralizablecertifiedunlearning].
Our approach is inspired by first-order methods such as [neel21a] [chien2024langevin] [chien2024stochastic]. In the aforementioned works, the learning algorithm $A$ is some variant of gradient descent on $f_{\mathcal{D}}$, and the unlearning algorithm $U$ is first initialized at the output weights of $A$ before proceeding with some number of additional gradient descent iterations on $f_{\mathcal{D^{\prime}}}$. Intuitively, because these works assume a unique minimizer or limiting distribution that attracts all gradient descent trajectories, the unlearning iterates will converge to the output $A(\mathcal{D^{\prime}})$. However, because gradient descent on nonconvex functions may converge to flat regions or saddle points, there is no guarantee that additional iterates after the learning process will make progress towards the destination of the retraining trajectory. Therefore, checkpointing $K$ iterates before the final timestep backtracks the algorithm to a steeper region of the loss landscape and draws the trajectory closer to the retraining trajectory.
3 Analyses
In the following theorem, we establish the unlearning and performance guarantees for our algorithm on nonconvex functions. For nonconvex functions, gradient descent might converge to local minima or saddle points, so we measure the performance by the average of the gradient norm over iterates. In differential privacy, gradient norm is a common performance metric for algorithms on nonconvex functions.
Theorem 3.1.
Let $\epsilon,\delta>0$ be fixed. Suppose for all $z\in\mathcal{Z}$, the loss function $f_{z}$ is $L$-Lipschitz smooth and the gradient is uniformly bounded by some constant $G$ so that $\lVert\nabla f_{z}(\theta)\rVert<G$ for all $\theta\in\Theta$. Let $D$ denote the original dataset of size $n$, let $Z\subset\mathcal{D}$ denote the unlearned dataset of size $m$, and let $\mathcal{D}^{\prime}=\mathcal{D}\backslash Z$ denote the retain set. Let the learning algorithm $A$ be initialized at $\theta_{0}$ and run for $T$ iterations with step-size $\eta\leq\min\{\frac{1}{L},\frac{n}{2(n-m)L}\}$. Let the standard deviation $\sigma$ of the Gaussian noise be defined as
$\sigma=\frac{2mG\cdot h(K)\sqrt{2\log(1.25/\delta)}}{Ln\epsilon},$ (2)
where $h(K)$ is a function that monotonically decreases to zero as $K$ increases from 00 to $T$ defined by
$h(K)=((1+\frac{\eta Ln}{n-m})^{T-K}-1)(1+\eta L)^{K}.$
Then $U$ is an $(\epsilon,\delta)$-unlearning algorithm for $A$ with noise $\sigma$. In addition,
$\frac{1}{T}\Big{[}\sum_{t=0}^{T-K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+\sum_{t=0}^{K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t})||^{2}+\mathbb{E}[||\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})||^{2}]\Big{]}\leq$ $\frac{2n(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K}))}{T\eta(n-m)}$ (3)
$+\frac{T-K-1}{T}(\frac{2G^{2}m}{n-m}+\frac{2L\eta Gm}{n-m})$
$\leq$ $O(\frac{n}{T(n-m)})+O(\frac{(T-K-1)m}{T(n-m)})$
where the expectation is taken with respect to the Gaussian noise added at the end of $U$, and where $O(\cdot)$ hides dependencies on $\eta$, $G$, and $L$.
Equation (3) states that the average of the gradient norm squared of the initial $T-K$ learning iterates, the $K-1$ unlearning iterates after, and the last perturbed iterate $\tilde{\theta}^{\prime\prime}$ decreases with increasing $T$ and $n$, showing that the algorithm converges to a stationary point with small gradient norm.
The proof of Theorem 3.1 is provided in Appendix B. The proof relies on carefully tracking the distance between the unlearning iterates and gradient descent iterates on $f_{\mathcal{D^{\prime}}}$. Like prior work in certified unlearning, our analysis relies on the Gaussian mechanism for differential privacy [dwork2014algorithmic], which implies that as long as the distance between the trajectories is bounded, we can add a sufficient amount of Gaussian noise to make the algorithm outputs ($\epsilon$, $\delta$)-indistinguishable. This allows us to determine the noise level $\sigma$ required to achieve unlearning. For the utility guarantees, we leverage the fact that gradient descent steps on $f_{\mathcal{D}}$ also make progress on $f_{\mathcal{D^{\prime}}}$ to obtain bounds that only depend on problem parameters and the initialization $\theta_{0}$ and do not depend on points along the training trajectory.
Theorem 3.1 underscores the “privacy-utility-complexity” tradeoff between our measure of unlearning, $\epsilon$, noise, $\sigma$, and the number of unlearning iterations, $K$. By construction, $K<T$, so our algorithm is more efficient than retraining. We can pick larger $K$ such that the noise required is arbitrarily small at the expense of computation, and when $K=T$ our algorithm amounts to a noiseless full retrain. Moreover, the standard deviation $\sigma$ inversely scales with the size of the dataset $n$, implying that unlearning on larger datasets require less noise. In contrast, [zhang2024towards] and [qiao2024efficientgeneralizablecertifiedunlearning] do not feature such data-dependent guarantees.
Remark.
The guarantee in Theorem 3.1 applies to unlearning a batch of $m$ data samples. Our algorithm also accommodates sequential unlearning requests. If, after unlearning $m$ points, an additional $k$ unlearning requests arrive, we can simply call the unlearning algorithm $M$ on the new retain set of size $n-m-k$. Notably, if the total number of unlearned data increases while $\sigma$ and $K$ stay constant, our unlearning guarantee worsens, which is consistent with other unlearning algorithms [zhang2024towards][chien2024langevin][Guo].
We can obtain faster convergence if we consider Polyak-Lojasiewicz (PL) functions, which are nonconvex functions that satisfy the following PL inequality.
Definition 3.2.
[karimiPL] For some function $f$, suppose it attains a global minimum value $f^{*}$. Then $f$ satisfies the PL inequality if for some $\mu>0$ and all $x$,
$\frac{1}{2}||f(x)||^{2}\geq\mu(f(x)-f^{*}).$ (4)
Since PL functions have a continuous basin of global minima, we can obtain empirical risk bounds. Corollary 3.3 states that the empirical risk converges linearly with both $T$ and $K$. The proof is provided in Appendix B.1.
Corollary 3.3.
Suppose the conditions of Theorem 3.1 hold and in addition, $f_{\mathcal{D}^{\prime}}$ satisfies the PL condition with parameter $\mu$. Let $f_{\mathcal{D^{\prime}}}^{*}$ represent the global optimal value of $f_{\mathcal{D^{\prime}}}$. Then
$\mathbb{E}[f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})]-f_{\mathcal{D^{\prime}}}^{*}\leq L\sqrt{d}\sigma+(1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}$
where $\sigma$ is defined in (2) and the expectation is taken with respect to the Gaussian noise added at the end of $U$.
Our performance guarantees demonstrate that more learning iterates $T$ correspond to better performance upon unlearning. For PL functions, the utility converges faster with increasing $T$ than for the general nonconvex case.
Although practical algorithms typically minimize the empirical risk, the ultimate goal of learning is to minimize the population risk, given by
$F(\theta)=\mathbb{E}_{z\sim\mathcal{Z}}[f_{z}(\theta)],$ (5)
in order to determine how well the model will generalize on unseen test data. We leverage results from [lei2021sharper] that relate the on-average stability bounds of algorithms on PL functions to their excess population risk.
Corollary 3.4.
Suppose the conditions of Theorem 3.1 hold and in addition, $f_{\mathcal{D}^{\prime}}$ satisfies the PL condition with parameter $\mu$. Let $F^{*}$ represent the global optimal value of $F$, defined in (5). Then
$\mathbb{E}[F(\tilde{\theta}^{\prime\prime})]-F^{*}\leq$ $L\sqrt{d}\sigma+\frac{2G^{2}}{(n-m)\mu}$
$+\frac{L}{2\mu}\left((1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}\right)$
where $\sigma$ is defined in (2) and the expectation is taken with respect to the i.i.d. sampling of $\mathcal{D}\sim\mathcal{Z}$ and the Gaussian noise added at the end of $U$.
Comparison with Prior Work.
To review, our algorithm is a first-order, black-box algorithm that provides $(\epsilon,\delta)$ unlearning while also maintaining performance and computational efficiency. We next compare our algorithm with the prior literature in this area. Earlier in Table 1, we review our algorithmic differences with existing convex and nonconvex certified unlearning algorithms. These algorithms all involve injecting some (Gaussian) noise to render the algorithm outputs probabilistically indistinguishable, whether it is added to the objective function, to the final model weights, or at each step of the training process. We are interested in comparing the noise-unlearning tradeoff of each of these algorithms.
Algorithm Standard Deviation of Injected Gaussian Noise
Newton Step [Guo] $\sigma=\frac{4LG^{2}\sqrt{2\log(1.5/\delta)}}{\lambda^{2}(n-1)\epsilon}$
Descent-to-delete (D2D) [neel21a] $\sigma=\frac{4\sqrt{2}M(\frac{L-\lambda}{L+\lambda})^{K}}{\lambda n(1-(\frac{L-\lambda}{L+\lambda})^{K})(\sqrt{\log(1/\delta)+\epsilon}-\sqrt{\log(1/\delta)})}$
Langevin Unlearning [chien2024langevin] No closed-form solution for $\sigma$ in terms of $\epsilon,\delta$
Constrained Newton Step [zhang2024towards] $\sigma=\frac{\left(\frac{2R(PR+\lambda)}{\lambda+\lambda_{min}}+\frac{32\sqrt{\log(d/\rho)}}{\lambda+\lambda_{min}}+\frac{1}{8}\right)LR\sqrt{2\log(1.25/\delta)}}{\epsilon}$
Hessian-free unlearning [qiao2024efficientgeneralizablecertifiedunlearning] $\sigma=2\eta G\frac{\sqrt{2\log(1.25/\delta)}}{\epsilon}\zeta_{T}^{-U}$
Our Work (R2D) $\sigma=\frac{2mG\cdot h(K)\sqrt{2\log(1.25/\delta)}}{Ln\epsilon}$
Table 2: Comparison of certified unlearning algorithms and their noise guarantees. For the other algorithms referenced in this table, $K$ is the number of unlearning iterates, $\lambda$ is the regularization constant or strongly convex parameter, $M$ is the Lipschitz continuity parameter, $d$ is the parameter dimension. We also have $||\theta||\leq R$ for algorithms that require uniformly bounded model weights. For [zhang2024towards], $P$ represents the Lipschitz constant of the Hessian, $\lambda_{min}$ represents the minimum eigenvalue of the Hessian, and $\rho$ represents a probability less than $1$. For [qiao2024efficientgeneralizablecertifiedunlearning], $\zeta_{T}^{-U}$ is a constant dependent on the upper and lower bounds of the Hessian spectrum that grows with the number of learning iterates $T$ for nonconvex objective functions. See Appendix A for an in-depth discussion of [chien2024langevin].
The last column of Table 2 lists the standard deviation $\sigma$ in terms of $\epsilon$, $\delta$, and other problem parameters for the other certified unlearning algorithms. Our result is analogous to that of [neel21a], due to algorithmic similarities such as weight perturbation at the end of training. However, the required noise in [neel21a] decays exponentially with unlearning iterations, while the noise decreases more slowly with increasing $K$ for our algorithm. This difference can be attributed to the strongly convex setting in [neel21a], which attracts trajectories to a global minimum.
Table 2 highlights some advantages our algorithm has over existing approaches. First, our theoretical guarantees do not require a uniform bound on the model weights $||\theta||\leq R$, nor does $\sigma$ depend on $R$, in contrast to [chien2024langevin] and [zhang2024towards]. Unlearning guarantees with strong dependence on $R$ are problematic, as illustrated in the following example. From the Gaussian mechanism of differential privacy, we have the following theorem.
Theorem 3.5.
[dwork2014algorithmic] Let $X\sim\mathcal{N}(\mu,\sigma^{2}\mathbb{I}_{d})$ and $Y\sim\mathcal{N}(\mu^{\prime},\sigma^{2}\mathbb{I}_{d})$. Suppose $\lVert\mu-\mu^{\prime}\lVert_{2}\leq\Delta$. Then for any $\delta>0$, $X$ and $Y$ are $(\epsilon,\delta)$-indistinguishable if
$\sigma=\frac{\Delta}{\epsilon}\sqrt{2\log(1.25/\delta)}.$
Suppose we have an arbitrary learning algorithm that outputs some model weights $\theta$ perturbed by Gaussian noise $\xi\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{d})$ and an arbitrary unlearning algorithm that also outputs some other model weights $\theta^{\prime}$ perturbed by Gaussian noise $\xi^{\prime}\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{d})$. Then we have
$\tilde{\theta}=\theta+\xi$
$\tilde{\theta^{\prime}}=\theta^{\prime}+\xi^{\prime}$
$||\theta-\theta^{\prime}||\leq 2R.$
Then by Theorem 3.5, our unlearning algorithm is $(\epsilon,\delta)$ unlearning as long as
$\sigma\geq\frac{\sqrt{2}R}{\sqrt{\log(1/\delta)+\epsilon}-\sqrt{\log(1/\delta)}}.$
In conclusion, an algorithm that outputs arbitrary weights in the feasible set can obtain unlearning with $\sigma=O(R)$, which is of the same order as the dependencies in [chien2024langevin] and [zhang2024towards].
Second, we note that [zhang2024towards] and [qiao2024efficientgeneralizablecertifiedunlearning] require a fixed injection of noise that does not vary with the number of data points or iterations, thereby implicitly upper bounding the utility of the resulting model. In contrast, our result clearly delineates the privacy-utility-efficiency tradeoff of our algorithm. Our noise also scales inversely with the size of the dataset $n$, such that unlearning on large datasets requires less perturbation. Finally, generally speaking, our algorithm requires fewer assumptions than prior works. We do not require Lipschitz continuity of the Hessian [zhang2024towards] or uniformly bounded Hessian eigenvalues [qiao2024efficientgeneralizablecertifiedunlearning] [zhang2024towards].
4 Conclusion
In this work, we introduce a novel certified unlearning algorithm for nonconvex functions that addresses the limitations of prior algorithms. We propose a first-order method that is black-box, in that it can be applied to a model pretrained with gradient descent, and does not have any dependence on the parameter set radius. In general, our algorithm is simple, easy to implement, and requires fewer assumptions than prior work.
A weakness of our algorithm is that it requires the storage of the model checkpoint for later unlearning. In addition, if the checkpoint was not saved during training, the full dataset is required to compute the checkpoint via the proximal point algorithm. However, once the checkpoint is saved, unlearning can occur with access to only the retained dataset.
Future directions of this research include adapting the algorithm to stochastic gradient descent, which is more commonly used for deep neural networks. However, extending the current analysis to the stochastic setting is challenging because for nonconvex functions, the noise at each step causes gradient descent trajectories to diverge exponentially from each other. We will also conduct experiments to empirically compare with other algorithms.
[…]
Appendix A Additional Discussion of Related Work
The work [chien2024langevin] proposes using noisy projected gradient descent to achieve certified unlearning, leveraging the existence and uniqueness of the limiting distribution of the training process as well as the boundedness of the projection set. Their guarantee shows that the privacy loss $\epsilon$ decays exponentially with the number of unlearning iterates.
However, $\sigma$, the noise added at each step, is defined implicitly with no closed-form solution. We can only assert that for a fixed number of iterations $K$, $\sigma$ must be at least $O(R)$ to obtain $\epsilon=O(1)$. Because $\sigma$ cannot be defined explicitly, it is difficult to implement this algorithm for a desired $\epsilon$. For example, when performing experiments for the strongly convex setting, which is a simpler mathematical expression, they require an additional subroutine to find the smallest $\sigma$ that satisfies the target $\epsilon$. As for the nonconvex setting, they state that “the non-convex unlearning bound… currently is not tight enough to be applied in practice due to its exponential dependence on various hyperparameters.”
Appendix B Proof of Theorem 3.1
Proof.
Let $Z$ be a dataset of $m$ data points we would like to unlearn, where $m<n$. Let $\mathcal{D}$ represent the original full dataset and $\mathcal{D^{\prime}}=\mathcal{D}\backslash Z$. Without loss of generality, define
$f_{\mathcal{D}}(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{z_{i}}(\theta)$
$f_{\mathcal{D}^{\prime}}(\theta)=\frac{1}{n-m}\sum_{i=1}^{n-m}f_{z_{i}}(\theta)$
such that we have
$f_{\mathcal{D}}=\frac{n-m}{n}f_{\mathcal{D}^{\prime}}(\theta)+\frac{1}{n}\sum_{i=n-m+1}^{n}f_{z_{i}}(\theta)$
$f_{\mathcal{D^{\prime}}}=\frac{n}{n-m}(f_{\mathcal{D}}(\theta)-\frac{1}{n}\sum_{i=n-m+1}^{n}f_{z_{i}}(\theta))$
Let $\{\theta_{t}\}_{t=0}^{T}$ represent the gradient descent iterates of the learning algorithm on $f_{\mathcal{D}}$, starting from $\theta_{0}$, and let $\{\theta^{\prime}_{t}\}_{t=0}^{T}$ be the iterates of the learning algorithm on $f_{\mathcal{D^{\prime}}}$ starting from the same $\theta_{0}$. Then we have
$\theta_{0}$ $=\theta^{\prime}_{0}$ (6)
$\theta_{t}$ $=\theta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})$ (7)
$\theta^{\prime}_{t}$ $=\theta^{\prime}_{t-1}-\eta\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime}_{t-1})$ (8)
Let $\{\theta^{\prime\prime}_{t}\}_{t=0}^{K}$ represent the gradient descent iterates of the unlearning algorithm starting at $\theta^{\prime\prime}_{0}=\theta_{T-K}$. Finally, let $\tilde{\theta}=\theta^{\prime\prime}_{K}+\xi$ denote the iterate with Gaussian noise added.
We first bound the distance between $\theta_{t}$ and $\theta^{\prime}_{t}$ as follows.
Lemma B.1.
Let $\{\theta_{t}\}_{t=0}^{T}$, $\{\theta^{\prime}_{t}\}_{t=0}^{T}$ be defined as in (6). Then
$\lVert\theta_{t}-\theta^{\prime}_{t}\rVert\leq\frac{2Gm}{Ln}[(1+\frac{\eta Ln}{n-m})^{t}-1]$
Proof.
We have
$\nabla f_{\mathcal{D^{\prime}}}(\theta)=\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta)-\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta))$
$\theta^{\prime}_{t}=\theta^{\prime}_{t-1}-\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1}))$
So we have for $\Delta_{t}=\theta_{t}-\theta^{\prime}_{t}$
$\Delta_{t}=\Delta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta)-\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1}))$
$=\Delta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1})$
$=\Delta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1}))-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1})$
$=\Delta_{t-1}+\eta\frac{m}{n-m}\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1}))-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}\nabla f_{z_{n}}(\theta^{\prime}_{t-1})$
We can take the absolute value of each side and by the triangle inequality
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{m}{n-m}||\nabla f_{\mathcal{D}}(\theta_{t-1})||+\eta\frac{n}{n-m}||\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1})||+\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}||\nabla f_{z_{n}}(\theta^{\prime}_{t-1})||$
Since the gradient on each data point is bounded by $G$, we have
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{m}{n-m}G+\eta\frac{n}{n-m}||\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1})||-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}G$
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{n}{n-m}||\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1})||+\frac{2\eta Gm}{n-m}$
By Lipschitz smoothness of the gradient, we have
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{n}{n-m}L||\theta^{\prime}_{t-1}-\theta_{t-1}||+\frac{2\eta Gm}{n-m}$
$||\Delta_{t}||\leq||\Delta_{t-1}||(1+\frac{\eta Ln}{n-m})+\frac{2\eta Gm}{n-m}$
Since we know that $||\Delta_{0}||=0$, evaluating this recursive relationship yields for $t>0$
$||\Delta_{t}||\leq\frac{2\eta Gm}{n-m}\sum_{i=0}^{t-1}(1+\frac{\eta Ln}{n-m})^{i}$
$||\Delta_{t}||\leq\frac{2\eta Gm}{n-m}\frac{(1+\frac{\eta Ln}{n-m})^{t}-1}{\frac{\eta Ln}{n-m}}$
$||\Delta_{t}||\leq 2Gm\frac{(1+\frac{\eta Ln}{n-m})^{t}-1}{Ln}$
This bound takes advantage of the difference between $f_{\mathcal{D}}$ and $f_{\mathcal{D^{\prime}}}$ as it decreases with large $n$, but it also grows exponentially with the number of iterates.
Lemma B.2.
Let $\{\theta^{\prime\prime}_{t}\}^{K}_{t=0}$ represent the gradient descent iterates on $f_{\mathcal{D^{\prime}}}$ starting at $\theta^{\prime\prime}_{0}=\theta_{T-K}$ such that
$\theta^{\prime\prime}_{t}=\theta^{\prime\prime}_{t-1}-\eta\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t-1})$
Then
$||\theta^{\prime}_{T}-\theta^{\prime\prime}_{K}||\leq||\theta_{T-K}-\theta^{\prime}_{T-K}||(1+\eta L)^{K}$
Proof.
Let $\Delta^{\prime}_{t}=\theta^{\prime}_{T-K+t}-\theta^{\prime\prime}_{t}$ such that $\Delta^{\prime}_{0}=||\theta^{\prime}_{T-K}-\theta^{\prime\prime}_{0}||$ and we bound it as follows
$\Delta^{\prime}_{t}=\theta^{\prime}_{T-K+t}-\theta^{\prime\prime}_{t}=\theta^{\prime}_{T-K+t-1}-\eta\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime}_{T-K+t-1})-\theta^{\prime\prime}_{t-1}+\eta\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t-1})$
$=\Delta^{\prime}_{t-1}-\eta\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime}_{T-K+t-1})+\eta\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t-1})$
$||\Delta^{\prime}_{t}||\leq||\Delta^{\prime}_{t-1}||+\eta||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime}_{T-K+t-1})-\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t-1})||$
By Lipschitz smoothness
$||\Delta^{\prime}_{t}||\leq||\Delta^{\prime}_{t-1}||+L\eta||\theta^{\prime}_{T-K+t-1}-\theta^{\prime\prime}_{t-1}||$
$||\Delta^{\prime}_{t}||\leq(1+\eta L)||\Delta^{\prime}_{t-1}||\leq||\Delta^{\prime}_{0}||(1+\eta L)^{t}$
Returning to the algorithm, suppose the learning algorithm has $T$ iterations and we backtrack for $K$ iterations. Then the difference between the output of the learning algorithm (without noise) on $f_{\mathcal{D^{\prime}}}$ and the unlearning algorithm would be
$||\theta^{\prime}_{T}-\theta^{\prime\prime}_{K}||\leq||\Delta_{T-K}||(1+\eta L)^{K}\leq\frac{2mG}{Ln}((1+\frac{\eta Ln}{n-m})^{T-K}-1)(1+\eta L)^{K}$
where the bound on the right hand side decreases monotonically as $K$ increases from 00 to $T$.
Let
$h(K)=((1+\frac{\eta Ln}{n-m})^{T-K}-1)(1+\eta L)^{K}$
Then the derivative is
$h^{\prime}(K)=(1+\eta L)^{K}[((1+\frac{\eta Ln}{n-m})^{T-K}-1)\log(1+\eta L)-(1+\frac{\eta Ln}{n-m})^{T-K}\log(1+\frac{\eta Ln}{n-m})]$
we observe that $h^{\prime}(K)<0$ for $K\in[0,T]$. Therefore $h(K)$ is decreasing.
Therefore, to achieve $\epsilon,\delta$-unlearning, we just need to set
$\sigma=\frac{||\theta^{\prime}_{T}-\theta^{\prime\prime}_{K}||\sqrt{2\log(1.25/\delta)}}{\epsilon}=\frac{2mG\cdot h(K)\sqrt{2\log(1.25/\delta)}}{Ln\epsilon}$
Now we move on to the utility guarantee. For nonconvex smooth functions, we know that for the gradient descent iterates $\theta^{\prime\prime}_{t}$, we have
$\frac{\eta}{2}\sum_{t=0}^{K}||\nabla f_{\mathcal{D}^{\prime}}(\theta^{\prime\prime}_{t})||\leq\sum_{t=0}^{K}f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t})-f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t+1})=f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{0})-f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K})$
Now we consider the progress of the iterates $\theta_{t}$ on $f_{\mathcal{D^{\prime}}}(\theta)$. By Lipschitz smoothness, we have
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})\leq f_{\mathcal{D^{\prime}}}(\theta_{t})+\langle\nabla f_{\mathcal{D^{\prime}}}(\theta_{t}),-\eta\nabla f_{\mathcal{D}}(\theta_{t})\rangle+\frac{L}{2}||\eta\nabla f_{\mathcal{D}}(\theta_{t})||^{2}$
We have
$\nabla f_{\mathcal{D}}(\theta_{t})=\frac{n-m}{n}\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})+\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta_{t})$
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})\leq f_{\mathcal{D^{\prime}}}(\theta_{t})-\eta\langle\nabla f_{\mathcal{D^{\prime}}}(\theta_{t}),\frac{n-m}{n}\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})+\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta_{t})\rangle+\frac{L\eta^{2}}{2}||\frac{n-m}{n}\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})+\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta_{t})||^{2}$
$\leq f_{\mathcal{D^{\prime}}}(\theta_{t})-\eta\frac{n-m}{n}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}-\eta\langle\nabla f_{\mathcal{D^{\prime}}}(\theta_{t}),\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta_{t})\rangle+\frac{L\eta^{2}}{2}||\frac{n-m}{n}\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})+\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta_{t})||^{2}$
$\leq f_{\mathcal{D^{\prime}}}(\theta_{t})-\eta\frac{n-m}{n}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}-\eta\langle\nabla f_{\mathcal{D^{\prime}}}(\theta_{t}),\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta_{t})\rangle+L\eta^{2}||\frac{n-m}{n}\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+L\eta^{2}||\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta_{t})||^{2}$
$\leq f_{\mathcal{D^{\prime}}}(\theta_{t})-\eta\frac{n-m}{n}(1-\frac{L\eta(n-m)}{n})||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}-\eta\langle\nabla f_{\mathcal{D^{\prime}}}(\theta_{t}),\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta_{t})\rangle+L\eta^{2}\frac{1}{n}\sum_{i=n-m+1}^{n}||\nabla f_{z_{i}}(\theta_{t})||^{2}$
$\leq f_{\mathcal{D^{\prime}}}(\theta_{t})-\eta\frac{n-m}{n}(1-\frac{L\eta(n-m)}{n})||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+\eta\frac{G^{2}m}{n}+L\eta^{2}\frac{mG}{n}$
Let the step-size $\eta$ be bounded such that $\eta\leq\frac{n}{2(n-m)L}$, then we have
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})\leq f_{\mathcal{D^{\prime}}}(\theta_{t})-\frac{\eta(n-m)}{2n}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+\frac{\eta G^{2}m}{n}+\frac{L\eta^{2}Gm}{n}$
Rearrange the terms to get
$\frac{\eta(n-m)}{2n}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}\leq f_{\mathcal{D^{\prime}}}(\theta_{t})-f_{\mathcal{D^{\prime}}}(\theta_{t+1})+\frac{\eta G^{2}m}{n}+\frac{L\eta^{2}Gm}{n}$
Sum from $t=0$ to $T-K-1$
$\frac{\eta(n-m)}{2n}\sum_{t=0}^{T-K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}\leq f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}(\theta_{T-K})+(T-K-1)(\frac{\eta G^{2}m}{n}+\frac{L\eta^{2}Gm}{n})$
$\sum_{t=0}^{T-K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}\leq\frac{2n}{\eta(n-m)}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}(\theta_{T-K})+(T-K-1)(\frac{\eta G^{2}m}{n}+\frac{L\eta^{2}Gm}{n}))$
$\sum_{t=0}^{T-K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}\leq\frac{2n}{\eta(n-m)}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}(\theta_{T-K}))+(T-K-1)(\frac{2G^{2}m}{n-m}+\frac{2L\eta Gm}{n-m})$
From standard analysis of gradient descent on nonconvex functions, we know
$\eta\sum_{t=0}^{K}||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t})||^{2}\leq f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{0})-f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K})$
$\sum_{t=0}^{K}||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t})||^{2}\leq\frac{1}{\eta}(f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{0})-f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K}))$
Summing the equations yields
$\sum_{t=0}^{T-K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+\sum_{t=0}^{K}||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t})||^{2}$
$\leq\frac{2n}{\eta(n-m)}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}(\theta_{T-K}))+(T-K-1)(\frac{2G^{2}m}{n-m}+\frac{2L\eta Gm}{n-m})+\frac{1}{\eta}(f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{0})-f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K}))$
$\leq\frac{2n}{\eta(n-m)}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K}))+(T-K-1)(\frac{2G^{2}m}{n-m}+\frac{2L\eta Gm}{n-m})$
We can expand $||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K})||^{2}$ as follows
$||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K})||^{2}$ $=||\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})||^{2}+||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K})||^{2}-||\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})||^{2}$
$=||\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})||^{2}+2\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})^{T}(\nabla f_{\mathcal{D^{\prime}}}(\theta_{K}^{\prime\prime})-\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime}))+||\nabla f_{\mathcal{D^{\prime}}}(\theta_{K}^{\prime\prime})-\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})||^{2}$
$=||\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})||^{2}+2\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})^{T}\xi+||\xi||^{2}$
So we have
$\frac{1}{T}\Big{[}\sum_{t=0}^{T-K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+\sum_{t=0}^{K-1}||\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{t})||^{2}+\mathbb{E}[||\nabla f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})||^{2}]\Big{]}\leq$ $\frac{2n(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}(\theta^{\prime\prime}_{K}))}{T\eta(n-m)}$
$+\frac{T-K-1}{T}(\frac{2G^{2}m}{n-m}+\frac{2L\eta Gm}{n-m})$
B.1 Proof of Corollary 3.3
As before, we first consider the gradient descent iterates on $f_{\mathcal{D}^{\prime}}$. For $\mu$-PL and smooth functions, we know that for the iterates $\theta^{\prime\prime}_{t}$, we have [karimiPL]
$f_{\mathcal{D}^{\prime}}(\theta^{\prime\prime}_{t})-f_{\mathcal{D}^{\prime}}^{*}\leq(1-\eta\mu)^{t}(f_{\mathcal{D}^{\prime}}(\theta^{\prime\prime}_{0})-f_{\mathcal{D}^{\prime}}^{*})$
Now we track the progress of the iterates $\theta_{t}$ on $f_{\mathcal{D^{\prime}}}$. By Lipschitz smoothness and the above analysis, we have
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})-f_{\mathcal{D^{\prime}}}(\theta_{t})\leq\langle\nabla f_{\mathcal{D^{\prime}}}(\theta_{t}),-\eta\nabla f_{\mathcal{D}}(\theta_{t})\rangle+\frac{L}{2}||\eta\nabla f_{\mathcal{D}}(\theta_{t})||^{2}$
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})-f_{\mathcal{D^{\prime}}}(\theta_{t})\leq-\eta\frac{n-m}{n}(1-\frac{L\eta(n-m)}{n})||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+\eta\frac{G^{2}m}{n}+L\eta^{2}\frac{mG}{n}$
Let the step-size $\eta$ be bounded such that $\eta\leq\frac{n}{2(n-m)L}$, then we have
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})-f_{\mathcal{D^{\prime}}}(\theta_{t})\leq-\frac{\eta(n-m)}{2n}||\nabla f_{\mathcal{D^{\prime}}}(\theta_{t})||^{2}+\frac{\eta G^{2}m}{n}+\frac{L\eta^{2}Gm}{n}$
By the PL inequality, we have
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})-f_{\mathcal{D^{\prime}}}(\theta_{t})\leq-\frac{\eta\mu(n-m)}{n}(f_{\mathcal{D^{\prime}}}(\theta_{t})-f_{\mathcal{D^{\prime}}}^{*})+\frac{\eta G^{2}m}{n}+\frac{L\eta^{2}Gm}{n}$
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})-f_{\mathcal{D^{\prime}}}(\theta_{t})\leq-\frac{\eta\mu(n-m)}{n}(f_{\mathcal{D^{\prime}}}(\theta_{t})-f_{\mathcal{D^{\prime}}}^{*})+\frac{\eta G^{2}m}{n}+\frac{L\eta^{2}Gm}{n}$
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})-f_{\mathcal{D^{\prime}}}^{*}\leq(1-\frac{\eta\mu(n-m)}{n})(f_{\mathcal{D^{\prime}}}(\theta_{t})-f_{\mathcal{D^{\prime}}}^{*})+\frac{\eta G^{2}m}{n}+\frac{L\eta^{2}Gm}{n}$
We evaluate this recursive relationship to obtain
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})-f_{\mathcal{D^{\prime}}}^{*}\leq(1-\frac{\eta\mu(n-m)}{n})^{t+1}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+\frac{1}{n}(\eta G^{2}m+L\eta^{2}Gm)\sum_{i=0}^{t}(1-\frac{\eta\mu(n-m)}{n})^{i}$
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})-f_{\mathcal{D^{\prime}}}^{*}\leq(1-\frac{\eta\mu(n-m)}{n})^{t+1}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+\frac{1}{n}(\eta G^{2}m+L\eta^{2}Gm)\sum_{i=0}^{t}(1-\frac{\eta\mu(n-m)}{n})^{i}$
$\leq(1-\frac{\eta\mu(n-m)}{n})^{t+1}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+\frac{1}{n}(\eta G^{2}m+L\eta^{2}Gm)\frac{n}{\eta\mu(n-m)}$
$f_{\mathcal{D^{\prime}}}(\theta_{t+1})-f_{\mathcal{D^{\prime}}}^{*}\leq(1-\frac{\eta\mu(n-m)}{n})^{t+1}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+\frac{G^{2}m+L\eta Gm}{\mu(n-m)}$
$f_{\mathcal{D^{\prime}}}(\theta_{0}^{\prime\prime})-f_{\mathcal{D^{\prime}}}^{*}=f_{\mathcal{D^{\prime}}}(\theta_{T-K})-f_{\mathcal{D^{\prime}}}^{*}\leq(1-\frac{\eta\mu(n-m)}{n})^{T-K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+\frac{G^{2}m+L\eta Gm}{\mu(n-m)}$
$f_{\mathcal{D^{\prime}}}(\theta_{K}^{\prime\prime})-f_{\mathcal{D^{\prime}}}^{*}\leq(1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}$ (9)
By Lipschitz smoothness, we have
$f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})-f_{\mathcal{D^{\prime}}}^{*}$ $\leq f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})-f_{\mathcal{D^{\prime}}}(\theta_{K}^{\prime\prime})+(1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}$ (10)
$\leq L||\tilde{\theta}^{\prime\prime}-\theta^{\prime\prime}_{K}||+(1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}$
$=L||\xi||+(1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}$
If we take the expectation on both sides with respect to the noise added at the end of the algorithm $U$, we have
$\mathbb{E}[f_{\mathcal{D^{\prime}}}(\tilde{\theta}^{\prime\prime})]-f_{\mathcal{D^{\prime}}}^{*}\leq L\sqrt{d}\sigma+(1-\frac{\eta\mu(n-m)}{n})^{T-K}(1-\eta\mu)^{K}(f_{\mathcal{D^{\prime}}}(\theta_{0})-f_{\mathcal{D^{\prime}}}^{*})+(1-\eta\mu)^{K}\frac{G^{2}m+L\eta Gm}{\mu(n-m)}$
B.2 Proof of Corollary 3.4
Bounds on generalization can be derived using seminal results in algorithmic stability [elisseeff05a]. Prior work on the generalization ability of unlearning algorithms focus on the strongly convex case [sekharirememberwhatyouwant] [qiao2024efficientgeneralizablecertifiedunlearning] [liu2023certified] [ullahTV], which feature excess risk bounds for the empirical risk minimizer when the loss is strongly convex [ShalevShwartz2009StochasticCO]. In contrast, algorithms on PL objective functions can at best satisfy pointwise [charles2017stability] or on-average [lei2021sharper] stability. The following result derives from Theorem 1 of [lei2021sharper]. We utilize the fact that for $w^{*}\in\arg\min_{\theta\in\Theta}F(\theta)$, we have $\mathbb{E}[f^{*}_{\mathcal{D}^{\prime}}]\leq\mathbb{E}[f_{\mathcal{D}^{\prime}}(w^{*})]=F(w^{*})=F^{*}$, and we also slightly modify the result to include the stronger bounded gradient assumption used in our work.
Lemma B.3.
[lei2021sharper] For $F$ defined in (5) and $\theta$ as the output of an algorithm dependent on $\mathcal{D^{\prime}}$, we have
$\mathbb{E}[F(\theta)-F^{*}]\leq\frac{2G^{2}}{(n-m)\mu}+\frac{L}{2\mu}\mathbb{E}[f_{\mathcal{D^{\prime}}}(\theta)-f^{*}_{\mathcal{D}^{\prime}}].$
We obtain the result by substituting in (9).